OpenStudy (anonymous):

What is the vertex, focus, and directix of the following? x^2=-3y Y^2=8x

OpenStudy (anonymous):

Hmm Well the directix for the first one is. 3/4

OpenStudy (anonymous):

And it seems the vertex is (0,0)

OpenStudy (mertsj):

Yes. It is concave downward so the directrix is a horizontal line. 4p = 3 so p = 3/4 so the focus is (0,-3/4) and the directrix is y = 3/4

OpenStudy (mertsj):

The second one opens to the right. It's vertex is also (0,0). This time 4p=8 so p = 2 so the focus is (2,0) and the directrix is the vertical line x = -2